# An $\ell^{1}$-Plug-and-Play Approach for Magnetic Particle Imaging Using a Zero Shot Denoiser with Validation on the 3D Open MPI Dataset

### v1

@@ 3.3 Experiment 1: Comparison on the OpenMPI Dataset
The OpenMPI Dataset [39] contains three phantoms designed to offer different challenges for the reconstruction: the shape phantom, the resolution phantom and the concentration phantom. We present both quantitative and qualitative results organized along the phantoms.
@@
Comparision Methods
@@
We display results for the proposed $\ell^{1}$-PnP approach and the DIP as well as the Tikhonov regularization method. For the comparison method of Tikhonov regularization, we have chosen its parameter $\lambda_{\mathrm{Tik}}$ optimally in the sense that we have performed a grid search and selected that value for which the $\mathrm{PSNR}_{\mathrm{max}}$ is maximal. The grid search proceeds as follows: we first perform reconstruction with parameters $10^{j}$ for $j=1,\cdots,10$ and consider $j^{\ast}$ the exponent that maximizes the PSNR; then we consider the values from $10^{j^{\ast}-1}$ to $10^{j^{\ast}+1}$ with step $10^{j^{\ast}-1}$ and again select the one maximizing PSNR. For the fine-tuning of the DIP comparison method, we have followed the approach proposed in [14]: we let the DIP run for 20000 iterations with three different gradient descent step sizes $\alpha_{i}=10^{-i}$ and we select the reconstruction corresponding to the parameter $\alpha_{i}$ and the iteration $s$ for which $\mathrm{PSNR}_{\mathrm{max}}$ is maximized (we refer to the Section S3 of the supplementary material for a detailed description of the architecture and the parameters we employ here). The optimal Tikhonov parameters and the iterations $s$ corresponding to the best $\mathrm{PSNR}_{\mathrm{max}}$ in the DIP are displayed in Tab. 1.
Parameters for the $\ell^{1}$-PnP Method
We choose the parameters $\mu_{k}$, $k\geq 1,$ $\lambda$ and $\alpha$ as well as the number of iterations $k$ as suggested in Sec 2.4. For $\mu_{0}$ however, we choose the optimal parameter $\lambda_{\mathrm{Tik}}$ computed for the comparison method above (cf. Tab. 1) instead of the estimate $\tilde{\lambda}_{\mathrm{Tik}}$ (cf. (14)). We note that $\lambda_{\mathrm{Tik}}$ is optimized w.r.t. PSNR involving the ground truth. We point out that we display the result upon convergence and do not consider an optimization on the iterations (as we do for the DIP).
Table 1: Concerning Experiment 1, we report the optimal Tikhonov parameters $\lambda_{\mathrm{Tik}},$ the number $s$ of the iteration with maximal PSNR for the DIP for the different gradient steps sizes $\alpha_{i}=10^{-i}$, $i=3,4,5,$ and the number $s$ of iterations of the $\ell^{1}$-PnP implied by the stopping criterion.
Phantom Method
Tikhonov DIP, $\alpha_{3}$ DIP, $\alpha_{4}$ DIP, $\alpha_{5}$ PnP
Shape Ph. $\lambda_{\mathrm{Tik.}}=23000$ s=110 s=250 s=1900 s=8
Res. Ph. $\lambda_{\mathrm{Tik.}}=8000$ s=4000 s=1900 s=3000 s=6
Conc. Ph. $\lambda_{\mathrm{Tik.}}=600$ s=4000 s=3000 s=15000 s=4
(a) Shape $\rho_{\text{GT}}$.
(b) Res. $\rho_{\text{GT}}$.
(c) Conc. $\rho_{\text{GT}}$.
(d) Shape $\rho_{\text{Tik}}$.
(e) Res. $\rho_{\text{Tik}}$.
(f) Conc. $\rho_{\text{Tik}}$.
(g) Shape $\rho_{\text{DIP}}$ with $\alpha_{3}$.
(h) Res. $\rho_{\text{DIP}}$ with $\alpha_{3}$.
(i) Conc. $\rho_{\text{DIP}}$ with $\alpha_{5}$.
(j) Shape $\rho_{\text{PnP}}$.
(k) Res. $\rho_{\text{PnP}}$.
(l) Conc. $\rho_{\text{PnP}}$.
Fig. 1: Selected slices for the Tikhonov reconstruction, the DIP with the $\alpha_{i}$ with the highest PSNR levels and the PnP (cf. Tab. 2). For the shape phantom we display the 10-th xz-slice, for the resolution phantom the 10-th xy-slice and for the concentration phantom the 13-th xy-slice.
@@ Results
Selected slices of the reconstructions are displayed in Fig. 1, we refer the interested reader to the supplementary material (Section S6) for a full display of all slices of the reconstructions.
Table 2: Results obtained with the $\ell^{1}$-PnP, the Tikhonov and DIP methods on the 3D Open MPI data set. We display the PSNR and SSIM scores for the reconstruction obtained maximizing with respect to PSNR for the Tikhonov method. For the DIP with different gradient steps sizes $\alpha=10^{-i}$, $i=3,4,5$ and the $\ell^{1}$-PnP we consider the reconstruction corresponding to the iteration for which the PSNR is maximal. In the last column, we display the results obtained with the $\ell^{1}$-PnP with parameter $\mu_{0}=\lambda_{\mathrm{Tik}}$ chosen as the optimal Tikhonov parameter and the other parameters according to Section 2.4.
Phantom Metric Method
Tikhonov DIP, $\alpha=10^{-3}$ DIP, $\alpha=10^{-4}$ DIP, $\alpha=10^{-5}$ $\ell^{1}$-PnP (our)
Shape $\mathrm{PSNR}$ 23.06 25.77 22.84 20.07 31.16
Phantom $\mathrm{SSIM}$ 0.505 0.852 0.668 0.508 0.951
Resolution $\mathrm{PSNR}$ 30.32 31.72 31.24 29.43 31.94
Phantom $\mathrm{SSIM}$ 0.505 0.702 0.606 0.432 0.730
Concentr. $\mathrm{PSNR}$ 36.54 37.74 37.65 38.3 37.87
Phantom $\mathrm{SSIM}$ 0.546 0.464 0.508 0.504 0.576
An overview of the results obtained in terms of $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ are displayed in Tab. 2. The $\ell^{1}$-PnP method obtains higher SSIM for all phantoms whereas the PSNR score of the $\ell^{1}$-PnP is significantly higher for the shape phantom, higher for the resolution phantom and lower but comparable for the concentration phantom.
Fig. 2: Example of a profile cut for the shape phantom. We consider the 9-th xz-slice and the cut along the 9-th x pixel.
Because the $19\times 19\times 19$ resolution of the reconstructions, the visual quality assessment of the differences between reconstructions can be difficult. For a better understanding of the differences we have plotted in Fig. 2 a 1D line plot that compares the graph of the reconstructions and the reference CAD model. In Fig. 2 we can see that the Tikhonov method smooths out the reference and takes negative values, whereas the the DIP tends to overshoot when compared to the PnP reconstruction. It is worth noting that the DIP can present artifacts in the reconstruction (cf. Fig. 0(h)).
@@
Finally, we have measured the wall clock reconstruction times of all methods. Because of the use of the SVD in this experiment, Tikhonov type reconstructions are very speedy on the GPU, in particular, one direct solution takes about $10^{-3}$ seconds. However, with increasing problem sizes in future MPI applications, a SVD might not always be available. For this reason we have run the experiments with $\ell^{1}$-PnP solving the linear systems with the Conjugated Gradient (CG) method in place of the direct solution with the SVD as well. For the CG method we have set a tolerance of $10^{-12}$ and 10000 maximum number of iterations.
@@
DIP 148.9 s 145.9 s 142.9 s
$\ell^{1}$-PnP (SVD) 1.6 s 1.35 s 0.86 s
$\ell^{1}$-PnP (CG) 2.5 s 1.44 s 1.24 s
Table 3: Runtime table. The runtime of the DIP (which is independent of the step size parameter $\alpha_{i}$) is compared with the runtime of the proposed $\ell^{1}$-PnP algorithm using both the SVD (rSVD) in the direct solution of the linear systems and the CG method.
From Tab. 3 we observe that 20000 iterations of the DIP take on average 146 seconds whereas the reconstruction performed with the $\ell^{1}$-PnP method takes between 1-3 seconds using the CG method and on average a bit more than a second using the SVD (rSVD) for the direct solution of the linear systems.
3.4 Experiment 2: Reconstructions with Decreasing Levels of Preprocessing for the $\ell^{1}$-PnP
In our second experiment we consider the shape phantom and study the proposed $\ell^{1}$-PnP algorithm with respect to different levels of preprocessing. In contrast to Experiment 1, the choice of all parameters is according to Sec 2.4, and thus does not involve any groundtruth. For computation and (hyper-)parameter estimation we rely on the SVD.
@@ 3.4 Experiment 2: Reconstructions with Decreasing Levels of Preprocessing for the $\ell^{1}$-PnP
We consider three different preprocessing levels for the shape phantom data. For all three preprocessing levels we consider the frequencies between 80 and 625 kHz. At first we consider whitening and a rSVD with reduced rank $K=2000$; secondly, we do not perform whitening but still consider the rSVD with $K=2000$; finally, we do not perform whitening and consider a full rank rSVD with $K=19^{3}=6859$.
@@
We choose the parameters according to Sec 2.4. (The smallest singular values $\sigma_{\mathrm{min}},$ the starting parameters $\mu_{0}$ as well as the resulting number of iterations are displayed in Tab. 4.)
Phantom Metric Pre-Processing
@@
Shape PSNR 30.54 29.34 31.19
Phantom SSIM 0.947 0.939 0.952
stop it. 9 11 10
$\sigma_{\mathrm{min}}$ 22.79 14875.63 6284.47
$\mu_{0}$ $51925.5$ $\sim 2.2\times 10^{10}$ $\sim 3.9\times 10^{9}$
Table 4: $\ell^{1}$- PnP reconstruction on the shape phantom using our parameter choice strategy (including stopping criterion); results with decreasing levels of preprocessing.

### v2

@@ 3.3 Experiment 1: Comparison on the OpenMPI Dataset
The OpenMPI Dataset [39] contains three phantoms designed to offer different challenges for the reconstruction: the shape phantom, the resolution phantom and the concentration phantom. We present both quantitative and qualitative results organized along the phantoms. For the validation of the parameters we have used the dataset described in Section 2.5.
@@
Parameter Validation
We display results for the proposed ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP approach, the DIP as well as the Tikhonov regularization method and the PP-MPI approach.
@@
For the comparison method of Tikhonov regularization, we have chosen its parameter $\lambda_{\mathrm{Tik}}$ optimally in the sense that we have validated it performing a grid search and selecting that value for which the average $\mathrm{PSNR}$ is maximal. The grid search proceeds as follows: we first perform reconstruction with parameters $10^{j}$ for $j=-6,-5,\dots,0,1,\cdots,18$ and consider $j^{\ast}$ the exponent that maximizes the PSNR; then we consider the values $k\cdot 10^{j^{\ast}-1}$ and $k\cdot 10^{j^{\ast}}$ with $k=1,\dots,9$ and again select the one maximizing the average PSNR.
For the fine-tuning of the DIP comparison method, we have followed the approach proposed in [14]: we let the DIP run for 20000 iterations with three different gradient descent step sizes $\alpha_{i}=10^{-i}$, $i=3,4,5$ on the validation dataset and we select the parameter $\alpha_{i}$ and the iteration $s$ for which the average $\mathrm{PSNR}$ is maximized. (We refer to Section 3.2 and [14] for details on the implementation.)
For the PP-MPI in  [4] we have validated the number of iterations and the variable $\epsilon$ in Equation (19). Consistently with [4], we consider $f=Au_{\mathrm{GT}}+\varepsilon\eta$ for $\eta\sim\mathcal{N}(0,\mathbb{I})$ with $\varepsilon=\sqrt{\lVert Au_{\mathrm{GT}}\rVert^{2}/(\mathrm{SNR}\cdot\lVert\eta\rVert^{2})}$, and $\mathrm{SNR}_{\mathrm{db}}=10\log_{10}(\mathrm{SNR})$. We performed 3000 total iterations with $\epsilon$ with $\mathrm{SNR}_{db}$ ranging from 20 to 30 with step size 1.
For both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP methods we have validated the parameter $\mu_{0}$ and the optimal number of iterations $n_{\mathrm{it}}$ again considering the optimal parameters that maximize the average $\mathrm{PSNR}$. For $\mu_{0}$ we have performed the same grid search as for the Tikhonov parameter.
Comparison Methods Proposed ZeroShot-PnP
Tikhonov DIP PP-MPI [4] Without $\ell^{1}$-Prior With $\ell^{1}$-Prior
Param. $\lambda_{\mathrm{Tik.}}=5\cdot 10^{5}$ $\alpha=10^{-3}$ $\mathrm{SNR}_{db}=25$ $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
$s=14000$ $n_{\mathrm{it}}=2170$ $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $25.28\pm 3.28$ $29.45\pm 3.07$ $26.10\pm 3.95$ $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.657\pm 0.140$ $0.646\pm 0.215$ $0.524\pm 0.195$ $0.813\pm 0.140$ $0.794\pm 0.187$
Table 1: Validation results on the hybrid dataset in Section 2.5. We observe that the proposed ZeroShot-PnP methods achieve the highest values w.r.t. both PSNR and SSIM on the hybrid datset.
An overview of the validated parameters is displayed in Table 1. We observe that on the validation dataset both the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP proposed in this paper achieve higher $\mathrm{PSNR}$ and $\mathrm{SSIM}$ values.
(a) Shape CAD reference.
(b) Shape Tikhonov.
(c) Shape DIP.
(d) Shape PP-MPI.
(e) Shape ZeroShot-PnP.
(f) Shape ZeroShot-$\ell^{1}$-PnP.
Figure 1: Reconstructions of the shape phantom. We display here the 10-th xz-slice.
(a) Res. CAD reference.
(b) Res. Tikhonov.
(c) Res. DIP.
(d) Res. PP-MPI.
(e) Res. ZeroShot-PnP .
(f) Res. ZeroShot-$\ell^{1}$-PnP.
Figure 2: Reconstructions of the resolution phantom. We display the 10-th xy-slice.
(a) Conc. CAD reference.
(b) Conc. Tikhonov.
(c) Conc. DIP.
(d) Conc. PP-MPI.
(e) Conc. ZeroShot-PnP.
(f) Conc. ZeroShot-$\ell^{1}$-PnP.
Figure 3: Reconstructions for the concentration phantom. We display the 13-th xy-slice. We observe that, even though the PSNR value of the DIP reconstruction is high (cf. Table 2), it is possible that parts of the phantom are missing.
@@ Results
Selected slices of the reconstructions of the three phantoms are displayed in Figures 12 and 3. The reconstructions where performed with the validated parameters in Table 1.
Phantom Metric Comparison Methods Proposed ZeroShot-PnP
Tikhonov DIP PP-MPI [4] with $\ell^{1}$-Prior without $\ell^{1}$-Prior
Shape $\mathrm{PSNR}_{\mathrm{max}}$ 22.64 28.96 21.14 31.84 31.87
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.489 0.871 0.769 0.954 0.954
Resolution $\mathrm{PSNR}_{\mathrm{max}}$ 30.64 32.89 30.09 32.1 32.09
Phantom $\mathrm{SSIM}_{\mathrm{max}}$ 0.593 0.737 0.519 0.731 0.730
Concentr. $\mathrm{PSNR}_{\mathrm{max}}$ 36.14 37.78 36.09 37.54 37.45
Phatom $\mathrm{SSIM}_{\mathrm{max}}$ 0.518 0.441 0.535 0.578 0.577
Table 2: Results obtained with the ZeroShot-$\ell^{1}$-PnP, the Tikhonov, the DIP and the PP-MPI methods on the 3D Open MPI dataset. We display the $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ scores for the reconstructions obtained using the validated parameter in Table 1.
An overview of the results obtained in terms of $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ are displayed in Tab. 2.
We can observe that the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP methods achieve similar results, and in some instances the additional $\ell^{1}$-prior can increase the PSNR score (cf. shape phantom). In any case, the reconstruction quality according to both PSNR and SSIM of ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP are higher than the baseline Tikhonov method and the previous PP-MPI approach. The reconstruction quality of the two methods proposed in this paper for MPI reconstruction are comparable in quality with the values obtained for the DIP. However, the DIP takes considerably more time to perform a reconstruction and involves a random initial state $z$ on which the quality of the reconstruction depends. Moreover, comparing the DIP reconstruction in Figure 3(c) with the CAD reference 3(a) and the ZeroShot-$\ell^{1}$-PnP reconstruction 3(f), it is possible that even if the PSNR value is the highest, some features of the phantom may not be reconstructed by the DIP.
Figure 4: Example of a profile cut for the shape phantom. We consider the 10-th xz-slice and the cut along the 9-th x pixel. We point out that the profiles obtained with the ZeroShot-PnP and the ZeroShot-$\ell^{1}$-PnP overlap almost perfectly.
Because the $19\times 19\times 19$ resolution of the reconstructions, the visual quality assessment of the differences between reconstructions can be difficult. For a better understanding of the differences we have plotted in Figure 4 a 1D line plot that compares the graph of the reconstructions and the reference CAD model. In Figure 4 we can see that the Tikhonov method smooths out the reference and takes negative values, whereas the the DIP and the PP-MPI tend to overshoot when compared to the ZeroShot-PnP reconstruction. The ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP lines are almost completely overlapping. We observe that the DIP can present artifacts in the reconstruction (cf. Figure 3(c) and 3(a)).
@@
Finally, we have measured the wall clock reconstruction times for all methods. Because of the use of the SVD in this experiment, Tikhonov type reconstructions are very speedy on the GPU, in particular, one direct solution takes about $6\cdot 10^{-4}$ seconds.
@@
Tikhonov $6.25\cdot 10^{-4}$ s $6.59\cdot 10^{-4}$ s $6.54\cdot 10^{-4}$ s
DIP 129.93 s 165.87 s 165.46 s
PP-MPI 190.08 s 183.24 s 184.47 s
ZeroShot-PnP 1.10 s 1.16 s 1.02 s
ZeroShot-$\ell^{1}$-PnP 0.72 s 0.74 s 0.74 s
Table 3: Wall clock runtime table of the examined methods. The times are computed by averaging 100 reconstructions on the Open MPI dataset. We observe the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP take the least amount of time for the reconstructions.
From Table 3 we observe that - using the validated parameters in Table 1 on the Open MPI Dataset - apart from the Tikhonov method, the ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP methods are the fastest in this comparison and achieve the results in Table 2 in less than 2 seconds.
3.4 Experiment 2: Study on the Differences with the PP-MPI Approach
In this experiment we study more closely the differences between the previous PP-MPI approach and the ZeroShot-PnP approaches proposed for MPI in this paper. In particular, we consider the PP-MPI algorithm with its denoiser trained on MPI-friendly data (PP-MPI-Denoiser) [4]; then, we consider the proposed Algorithm 1 first with the PP-MPI-Denoiser and then with the ZeroShot-Denoiser [70]. The objective here is to show that the employment of the zero-shot denoiser achieves competitive results when compared with the previous PPP-MPI-Denoiser. Moreover, this experiment shows that with the reconstruction parameter automatically selected as proposed in Section 2.4, it is possible to enhance the reconstruction results also if the PP-MPI-Denoiser is employed.
@@ 3.4 Experiment 2: Study on the Differences with the PP-MPI Approach
We consider the same preprocessing as in Experiment 1.
Parameter Validation
We choose the parameters according to Section 2.4. The validated starting parameters $\mu_{0}$ as well as the resulting number of iterations are displayed in Table 4.
PP-MPI-Denoiser [4]
PP-MPI HQS-Splitting Eq.(12) $\ell^{1}$-HQS-Splitting Eq. (9)-(11)
Param. $\mathrm{SNR}_{db}=25$ $\mu_{0}=4\cdot 10^{11}$ $\mu_{0}=6\cdot 10^{16}$
Iteration $n_{\mathrm{it}}=2170$ $n_{\mathrm{it}}=9$ $n_{\mathrm{it}}=4$
PSNR $26.10\pm 3.95$ $27.75\pm 5.20$ $27.68\pm 5.15$
SSIM $0.524\pm 0.195$ $0.567\pm 0.224$ $0.644\pm 0.215$
Proposed Denoiser [70]
ZeroShot-PnP ZeroShot-$\ell^{1}$-PnP
Param. $\mu_{0}=2\cdot 10^{5}$ $\mu_{0}=3\cdot 10^{5}$
Iteration $n_{\mathrm{it}}=7$ $n_{\mathrm{it}}=5$
PSNR $30.45\pm 5.04$ $29.77\pm 5.82$
SSIM $0.813\pm 0.140$ $0.794\pm 0.187$
Table 4: Validation results on the hybrid dataset in Section 2.5. We observe that the employment of Algorithm 1 in combination with the PP-MPI-Denoiser leads to a lower number of iterations and to increased PSNR and SSIM scores when compared with the PP-MPI algorithm. However, the highest PSNR and SSIM score are obtained using the proposed ZeroShot-Denoiser (which has not been trained on MPI data).
Results
From the PSNR and SSIM values obtained on the hybrid dataset in Table 4 we can observe that both the HQS-Splitting (Equation (12)) and the $\ell^{1}$-HQS-Splitting of Equations (9)-(11) with the PP-MPI-Denoiser achieve higher metric scores than the original ADMM-based PP-MPI algorithm. Moreover, these improvements are achieved in less than 10 iterations in both cases. We observe that in the algorithm with the PP-MPI-Denoiser, the HQS-Splitting algorithm achieves the highest PSNR value whereas the $\ell^{1}$-HQS-Splitting achieves the highest SSIM score without compromising too much the PSNR: this means that there are cases in which one might prefer one method over the other depending on the objective. However, from the results in Table 4 it is also clear that the combination of the proposed splittings and the proposed denoiser [70], i.e., the ZeroShot-PnP and the ZeroShot-$\ell^{1}$-PnP approaches, yields the results with the highest PSNR and SSIM scores.
3.5 Experiment 3: Reconstructions with Decreasing Levels of Preprocessing for the ZeroShot-$\ell^{1}$-PnP
In our third experiment we consider the shape phantom and study the proposed ZeroShot-PnP and ZeroShot-$\ell^{1}$-PnP algorithms with respect to different levels of preprocessing.
Preprocessing
We consider three different preprocessing levels for the shape phantom data. For all three preprocessing levels we consider the frequencies between 80 and 625 kHz. At first we consider whitening and an rSVD with reduced rank $K=2000$; secondly, we do not perform whitening but still consider the rSVD with $K=2000$; finally, we do not perform whitening and consider a full rank rSVD with $K=19^{3}=6859$.
@@ 3.5 Experiment 3: Reconstructions with Decreasing Levels of Preprocessing for the ZeroShot-$\ell^{1}$-PnP
Consistently with Experiment 1, we validate the parameters using our hybrid dataset. The validated parameters are displayed in Table 5.
Method Quantity Pre-Processing
@@
ZeroShot PSNR 31.84 31.05 31.30
PnP SSIM 0.954 0.946 0.950
$n_{\mathrm{it}}$ 6 19 8
$\mu_{0}$ $2\cdot 10^{5}$ $6\cdot 10^{16}$ $10^{9}$
ZeroShot PSNR 31.87 31.14 29.8
$\ell^{1}$-PnP SSIM 0.954 0.947 0.956
$n_{\mathrm{it}}$ 4 3 74
$\mu_{0}$ $3\cdot 10^{5}$ $7\cdot 10^{8}$ $5\cdot 10^{10}$
Table 5: The ZeroShot-$\ell^{1}$-PnP adn ZeroShot-PnP validated parameters and metric computed on the shape phantom with different levels of preprocessing.